OpenStudy (anonymous):

prove that 2+ root5 is irrational number

OpenStudy (anonymous):

WELCOME TO OPEN STUDY and bc it haas many numbers after the decimal point in a random order here is the proof 2+root5=4.2360679774997896964091736687313

geerky42 (geerky42):

Here's the better structured proof: We will prove it using proof by contradiction. Suppose \(2+\sqrt5\) is rational \(r\), so that \[2+\sqrt5=r\]Squaring both sides, we have \(4+4\sqrt5+5=r^2\), which means that \(\sqrt{5}=\dfrac{r^2-9}{4}\) Since the set of rational numbers is closed under division, multiplication, and subtraction, \(\dfrac{r^2-9}{4}\) is rational. However it is known that \(\sqrt5\) is irrational. We have a contradiction. Therefore \(2+\sqrt5\) is irrational. \(\blacksquare\)

OpenStudy (anonymous):

Another way to see \(2+\sqrt{5}\) is irrational is as follows. If \(x=2+\sqrt{5}\), then: $$x=2+\sqrt{5}\Longrightarrow x-2=\sqrt{5}\Longrightarrow \left(x-2\right)^2=5$$ $$\Longrightarrow x^2-4x+4=5\Longrightarrow x^2-4x-1=0.$$Thus \(2+\sqrt{5}\) is a root of the polynomial \(x^2-4x-1\). If \(2+\sqrt{5}\) were rational, it would be a rational root of the the above polynomial. However, using the rational roots theorem, we see the only possible rational roots are \(\pm 1\), neither of which are actually roots. Thus \(2+\sqrt{5}\) cannot be rational.